Chris Company uses the perpetual inventory system and had the following activity during the current monthly period.

November 1:             Beginning Inventory of 100 units @$20

November 5:               Purchased 100 units @ $22

November 8:               Purchased 50 units @$23

November 16:            Sold 200 units @ $45

November 19:             Purchased 50 units @ $25

Using the weighted-average inventory method, the company's ending inventory would be reported at:

A)   $2,200.

B) $2,320.

C) $2,250.

D) $2,400.

E) $2,270.
Answer #1

Answers

  • Units available for Sale = 100 + 100 + 50 + 50 = 300 units
  • Units sold = 200
  • Ending Inventory = 300 – 200 = 100 units
  • Workings

Average Method

Cost of Goods available for sale

Units

Cost/unit

COG for sale

Beginning Inventory

100

$                20.00

$                        2,000.00

Purchases:

Nov 5

100

$                22.00

$                        2,200.00

Nov 8

50

$                23.00

$                        1,150.00

Nov 19

50

$                25.00

$                        1,250.00

TOTAL

300

$              22.000

$                        6,600.00

  • Weighted average cost per unit on Nov 16 = ($2000 + 2200 + 1150) / (100 + 100 + 50) = $ 21.40
  • Cost of Goods Sold = 200 units x $ 21.40 = $ 4280
  • Ending Inventory= Cost of Goods available for sale – Cost of Goods Sold
    = $ 6600 - $ 4280
    = $ 2,320
  • Correct Answer = Option ‘B’ $ 2,320